### v2

# Regular AdS3 black holes from regularized Gauss-Bonnet coupling

Abstract
We obtain a three-dimensional bi-vector-tensor theory of the generalized Proca class by regularizing the Gauss-Bonnet invariant within the Weyl geometry. We show that the theory admits a regular AdS3 black hole solution with primary hairs. Introducing a deformation in the theory, a different regular AdS3 black hole solution is obtained. Charged generalizations of these solutions are given by coupling to Born-Infeld electrodynamics.
1 Introduction
Thanks to its relative simplicity, gravity in three dimensions (3d) serves as a useful theoretical laboratory. For example, 3d general relativity (GR) with a negative cosmological constant admits a black hole solution named as the Banados-Teitelboim-Zanelli (BTZ) black hole [1, 2], which is locally isomorphic to the AdS3 spacetime but still has all the important properties of a black hole. The geometry is characterized by the mass and the angular momentum, making it globally different than the AdS3 spacetime. It also has a well-defined event horizon and laws of black hole thermodynamics are satisfied. One of the most important challenges of a quantum theory of gravity is to provide a microscopic derivation of the black hole entropy. One motivation to study gravity in 3d is that it can be derived without any direct reference to string theory or supersymmetry [3] by employing AdS3/CFT2 correspondence (see [4] for a review).
Another important problem in gravitational physics is that there is no way to avoid singularities at the center of black holes or in the early universe within the standard framework of general relativity [5, 6, 7]. This leads to the divergence of certain physical observables such as tidal forces, and therefore looking for some mechanisms solving the problem within GR or trying to modify the theory such that the resulting solutions have no singularities are needed. In the light of the success achieved in the microscopic origin of the black hole entropy, one might try to study this singularity problem in 3d to gain insight into the problem.
The most natural approach to deal with the singularity problem is to try to obtain a regular black hole solution of GR by assuming a perfect fluid as a source (see [8] for such an investigation in 3d). Although it is possible to calculate the global charges and study the thermodynamics of solutions obtained with this procedure, there is not much information available about a potential mechanism or a modification of GR that would resolve the singularities. Another possible method is coupling GR to non-linear electrodynamics and successful examples can be found in [9, 10, 11].
Taking an effective field theory perspective, GR should only be seen as the leading term of an infinite series where the higher order terms are probed at high energies. Therefore, it is particularly important to obtain regular black hole solutions from theories with higher curvature corrections. The first 3d example was obtained in [12] (see [13] for the charged generalization), where a particular scalar-tensor theory, termed as the generalized electromagnetic quasi-topological gravity, admits regular black hole solutions provided that the scalar field is chosen to be linear in the azimuthal angle.
In this work, building upon our earlier work [14], we employ a bi-vector-tensor coupling obtained by a regularization of the Gauss-Bonnet (GB) invariant within the Weyl geometry and 3d find regular black hole solutions. The GB invariant is a topological term in 4d and identically vanishes in 3d. However, as shown in [15], a well-defined coupling that contributes to field equations in 4d can be obtained, and it gives rise to black hole solutions with a primary hair, i.e. an additional integration constant that modifies the geometry. By using the regularized GB coupling with two Weyl vectors, regular black hole solutions in 4d were obtained in [16]. In [14], we presented AdS3 black holes with primary hair by employing the 3d version of the regularized coupling. Now, in complete parallelism with the 4d story, we obtain regular AdS3 black hole solutions by considering two Weyl vectors.
The outline of this paper is as follows: In Section 2, we discuss the regularization of the GB invariant with one and two Weyl vectors. In Section 3, in order to introduce our methodology and notation, we revisit the solution arising from the regularized GB coupling with a single Weyl vector as given in [14]. In Section 4, we present the regularization of the GB invariant with two Weyl vectors. Then, we give regular AdS3 black hole solutions, and charged versions by coupling the theory to Born–Infeld electrodynamics. We present our conclusions in Section 5.
2 3d regularized Gauss–Bonnet invariant with two Weyl vectors
Lovelock gravity is a higher-curvature generalization of GR that shares many important physical properties of it (see [17] for a review). The lowest order correction that is quadratic in curvature, the GB invariant is given by
$\mathcal{G}=R^{2}-4R_{\mu\nu}R^{\mu\nu}+R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}.$ (2.1)
As mentioned in the Introduction, it contributes to field equations only when $d>4$. However, it was realized that there exists regularization mechanisms that allow us to obtain well defined Lagrangians giving rise to a nontrivial dynamical contribution in the lower dimensions. The first example yields scalar-tensor theories of Hordenski class [18, 19] and the resulting theories have quite interesting properties (see [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]). Recently, another regularization scheme within the Weyl geometry generating vector-tensor theories of the generalized Proca class was discovered [39]. Starting from a torsionless Weyl correction $\widetilde{\Gamma}^{\alpha}_{\,\mu\nu}$ with the corresponding covariant derivative $\widetilde{\nabla}_{\mu}$, the Weyl vector $W$ is defined by the following relation
$\widetilde{\nabla}_{\alpha}g_{\mu\nu}=-2g_{\mu\nu}W_{\alpha},$ (2.2)
from which the Weyl connection can easily be found in terms of the Levi–Civita connection and the Weyl vector (see reference [40, 41, 42, 43] for details). When the GB term is calculated with the Weyl connection, it reads
$\widetilde{\mathcal{G}}[W]=\mathcal{G}+(d-3)\nabla_{\mu}J^{\mu}+(d-3)(d-4)\mathcal{L},$ (2.3)
where
$J^{\mu}$ $=8G^{\mu\nu}W_{\nu}+4(d-2)\left[W^{\mu}\left(W^{2}+\nabla_{\nu}W^{\nu}\right)-W_{\nu}\nabla^{\nu}W^{\mu}\right],$ (2.4)
$\mathcal{L}$ $=4G^{\mu\nu}W_{\mu}W_{\nu}+(d-2)\left[4W^{2}\nabla_{\mu}W^{\mu}+(d-1)W^{4}\right],$ (2.5)
with $W^{2}=W_{\mu}W^{\mu}$, $W^{4}=(W^{2})^{2}$, and $\mathcal{G}$ is the GB invariant calculated with the Levi–Civita connection. Disregarding the boundary terms, the regularized GB invariant with a single Weyl vector is defined as
$\mathcal{G}[W]=\lim_{d\to p}\frac{\widetilde{\mathcal{G}}[W]-\mathcal{G}}{d-p},\quad\quad p=3,4.$ (2.6)
The two GB terms $\widetilde{\mathcal{G}}[W]$ and $\mathcal{G}$, and therefore their difference, do not contribute to the field equations. However, dividing the difference by dpd-p, which corresponds to scaling the relevant coupling constant, and then setting $d=p$ give a Lagrangian which yields non-trivial dynamics in $d=p$. One can repeat the same exercise by taking the difference between two GB invariants with different Weyl vectors as follows [16]
$\lim_{d\to p}\frac{\widetilde{\mathcal{G}}[B]-\widetilde{\mathcal{G}}[C]}{d-p}=\mathcal{G}[B]-\mathcal{G}[C].$ (2.7)
As a result, one obtains the difference of the regularized GB invariant for two different Weyl vectors. This coupling, when supplemented by the Einstein-Hilbert term, was shown to admit regular black holes in 4d. Here, we will use the 3d version of this coupling to obtain regular AdS3 black holes. It reads [14]
$\mathcal{G}[W]=-4G^{\mu\nu}W_{\mu}W_{\nu}-4W^{2}\nabla_{\mu}W^{\mu}-2W^{4},$ (2.8)
for a Weyl vector $W$.
Before we attempt to find the black hole solutions of the resulting by bi-vector-tensor theory, as a warm-up exercise, we will revisit the solution of the theory with the single Weyl vector given in [14] and explain the shortcut used in [44] recently.
3 Solution in [14] revisited
We define the regularized vector-tensor Einstein–Gauss–Bonnet theory by the following action
[…]
where $\Lambda_{0}$ is the bare cosmological constant, $\mathcal{G}[W]$ is the regularized GB invariant with the Weyl vector $W$ given in (2.8) and $\beta$ is the corresponding coupling constant. By assuming a line element of the form
[…]
and the following vector components
[…]
$L_{\text{red}}$ $=L^{\text{EH}}_{\text{red}}+\beta L_{\text{red}}^{[W]},$ (3.4)
$L^{\text{EH}}_{\text{red}}$ $=N(\Lambda_{0}r+f^{\prime}/2),$ (3.5)
$L_{\text{red}}^{[W]}$ $=\frac{rw_{0}^{4}}{f^{2}N^{3}}-N^{\prime}\left[\frac{2rw_{0}^{2}w_{1}}{N^{2}}+2f^{2}w_{1}^{2}(1+rw_{1})\right]-\frac{w_{0}^{2}\left[(1+2rw_{1})f^{\prime}+2f(w_{1}+rw_{1}^{2}+rw_{1}^{\prime})\right]}{fN}$
$+N\left[fw_{1}^{2}(1+2rw_{1})f^{\prime}+f^{2}w_{1}^{2}(2w_{1}+rw_{1}^{2}+2rw_{1}^{\prime})\right].$ (3.6)
Here, $L^{\text{EH}}_{\text{red}}$ is the contribution of the Einstein-Hilbert term (with the bare cosmological constant $\Lambda_{0}$) and $L_{\text{red}}^{[W]}$ is the contribution of the regularized GB invariant with the Weyl vector $W$. Equations of motion for the functions $(N,f,w_{0},w_{1})$ can be found by extremizing the action. However, as first observed in [15] and applied to the 3d theory in [14], in order to decouple the equations, it is necessary to realize that the equations are invariant under the global symmetry
$w_{0}^{2}\to w_{0}^{2}+2\lambda fN(1+rw_{1})+\lambda^{2}r^{2},\quad\quad w_{1}\to w_{1}+\lambda\frac{r}{Nf},$ (3.7)
from which the following Noether charge follows
$Q=\frac{r^{2}w_{0}^{2}}{fN^{2}}-(1+rw_{1})^{2}f.$ (3.8)
Using this Noether charge together with the consistency of equations of motion, one finds that $N^{\prime}=0$ and therefore, one may take $N=1$. The metric function is found as
$f_{\pm}=-(m+q)+\frac{r^{2}}{2\beta}\left(1\pm\sqrt{1+4\beta\Lambda_{0}+\frac{4\beta q}{r^{2}}}\,\right).$ (3.9)
We see that an additional integration constant $q$ modifies the geometry, i.e., we have a black hole solution with a primary hair. The vector components can be found in terms of the metric function $f$, and integration constants $(m,q,c)$, where $c$ is another integration constant that does not modify the geometry. The norm of the Weyl vector is related to this constant on-shell, i.e. $W^{2}=-2c$. It turns out that the global symmetry (3.7) with $\lambda=-c$ kills the effect of the constant $c$ on the vector components, and one obtains a solution with a Weyl vector whose norm vanishes on-shell ($W^{2}=0$).
The solutions of the theory in general $d$-dimensions are studied in [44]. It was observed that, since the ultimate aim is to find the metric function and the norm of the vector field can always be made null by the global symmetry (3.7), one can use the following steps to find the solution in a theory containing the vector-tensor coupling in (2.8):
  1. 1.
    Find the reduced action by using the ansatz (3.2, 3.3) and derive the equations of motion.
  2. 2.
    Take
    $w_{1}=\frac{w_{0}}{Nf},$ (3.10)
    with which the norm of the Weyl vector vanishes ($W^{2}=0$). Thanks to this choice, one can easily get the solution where there is no additional integration constant (other than the hair parameter) appear in the vector components. This is equivalent to setting the additional constant to zero at the end by using the global symmetry of the reduced action, as explained before.
  3. 3.
    Then, one sees that the consistency of equations of motion implies $N=1$ and they now become decoupled. Solving $w_{0}$ first, the metric function $f$ can be easily found.
In the next section, we will apply this procedure to the bi-vector-tensor theory and its deformed version.
4 Regular black hole solutions
Now, we are prepared to study the solutions of the theory defined by the action
[…]
where $\ell$ is a constant of length dimension. For the Weyl vectors $B$ and $C$, similar to (3.3), we take
[…]
The contributions $L_{\text{red}}^{[B]}$ and $L_{\text{red}}^{[C]}$ to the reduced Lagrangian can be found by the replacement $W\rightarrow B,C$ in (3.6).
In 4d, by introducing a deformation, a different black hole solution was found [16]. Therefore, we do the same and consider also the following action with the deformation parameter $\kappa$
[…]
for which the reduced Lagrangian is given by
$L_{\text{red}}=L^{\text{EH}}_{\text{red}}+\ell^{2}\left(L_{\text{red}}^{[B]}-\kappa L_{\text{red}}^{[C]}\right).$ (4.4)
Note that this deformed coupling follows from the regularization $\lim_{d\to 3}\frac{\widetilde{\mathcal{G}}[B]-\kappa\widetilde{\mathcal{G}}[C]}{d-3}=\mathcal{G}[B]-\kappa\,\mathcal{G}[C]$.
By the method of [44] that we have outlined in the previous section, studying the solutions of the undeformed ($\kappa=1$) and the deformed ($\kappa\neq 1$) theory is a straightforward exercise. We now present regular black hole solutions and their charged versions by coupling the theories to Born–Infeld electrodynamics. In Appendix A, charged black hole solutions obtained by coupling to Maxwell’s electrodynamics are given. Due to the logarithmic form of the Coulomb potential in 3d, they possess a curvature singularity.
4.1 Undeformed theory ($\kappa=1$)
As mentioned before, after obtaining the equations of motion by extremizing reduced action in (4.4), analogous to (3.10), we choose
$b_{1}=\frac{b_{0}}{Nf},\quad\quad\quad\quad c_{1}=\frac{c_{0}}{Nf},$ (4.5)
Then, the consistency of the equations forces $N=1$ and one finds
$b_{0}=\frac{q_{b}}{r}-\frac{f}{2r},\quad\quad c_{0}=\frac{q_{c}}{r}-\frac{f}{2r},$ (4.6)
where $q_{b}$ and $q_{c}$ are the (constant) hair parameters. By the choice (4.5) and the condition $N=1$, the Weyl vectors $B$ and $C$ are null on-shell. With these vector components, the metric function can be found to be
$f=\frac{r^{2}(-\Lambda_{0}r^{2}-m)+4\ell^{2}(q_{b}^{2}-q_{c}^{2})}{r^{2}+4\ell^{2}(q_{b}-q_{c})},$ (4.7)
where we have introduced another integration constant $m$. As $r\to\infty$, the metric function behaves as
$f=-\Lambda_{0}r^{2}-m+4\ell^{2}(q_{b}-q_{c})\Lambda_{0}+\mathcal{O}(1/r^{2}),$ (4.8)
which shows that the effective cosmological constant is equal to the bare cosmological constant introduced in the action ($\Lambda=\Lambda_{0}$). Later, we will see that this is not case for the solutions of the deformed theory.
As long as $q_{b}>q_{c}$, any singularity at a positive $r$ is avoided. Near the origin $r\to 0$, one finds
$R=\frac{3(m+q_{b}+q_{c})}{2\ell^{2}(q_{b}-q_{c})}+\mathcal{O}(r^{2}),$ (4.9)
and the only other independent curvature invariants in 3d ($R^{\mu}_{\ \nu}R^{\nu}_{\ \mu}$ and $R^{\mu}_{\ \rho}R^{\rho}_{\ \nu}R^{\nu}_{\ \mu}$) [45, 46, 47] are also finite at the origin. Therefore, we conclude that it is a regular black hole solution with two primary hairs. Note that $q_{b}=q_{c}$ makes $B=C$ on-shell, and one obtains a stealth BTZ black hole solution.
4.2 Undeformed theory ($\kappa=1$) coupled to Born–Infeld electrodynamics
We now couple the theory to Born–Infeld electrodynamics with the Lagrangian [48]
$\mathcal{L}_{\text{BI}}=b^{2}\left(1-\sqrt{1+\frac{F_{\mu\nu}F^{\mu\nu}}{2b^{2}}}\,\right),$ (4.10)
where $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the field strength tensor of the $U(1)$ gauge field $A_{\mu}$. In the $b\to\infty$ limit of the Born-Infeld parameter, it reduces to the Lagrangian of Maxwell’s theory in (A.1). By assuming
[…]
the scalar potential solving the equation of motion can be taken as [49]
$\phi=-q_{e}\log\left(\frac{r+\sqrt{r^{2}+\frac{q_{e}^{2}}{b^{2}}}}{2}\right).$ (4.12)
Using the vector components in (4.5, 4.6) together with the contribution from the Born–Infeld electrodynamics, the metric function is found as
$f=\frac{r^{2}[-\Lambda_{0}r^{2}-m+\frac{q_{e}^{2}}{2}(\frac{1}{1+\psi}-\text{arctanh}(\psi))]+4\ell^{2}(q_{b}^{2}-q_{c}^{2})(\frac{b^{2}r^{2}}{q_{e}^{2}}+1)}{r^{2}+4\ell^{2}(q_{b}-q_{c})}$ (4.13)
where
$\psi=\frac{r}{\sqrt{r^{2}+q_{e}^{2}/b^{2}}}.$ (4.14)
The effective cosmological constant is again equal to the bare one ($\Lambda=\Lambda_{0}$). Near the origin, the Ricci scalar behaves as
$R=\frac{3\left(2q_{b}^{2}+m\sqrt{\kappa}\,\ell^{2}\left|(\kappa-1)q_{b}\right|\right)}{2(\kappa-1)\ell^{2}q_{b}^{2}}+\mathcal{O}(r^{2}),$ (4.15)
and the other curvature invariants are well-defined at the origin, which show that it is a regular black hole as long as $q_{b}>q_{c}$. When $q_{b}=q_{c}$, we have the stealth black hole solution of Einstein–Born–Infeld theory [48].
4.3 Deformed theory ($\kappa\neq 1$)
The deformed theory can also be solved through the same procedure and the vector components take the same form in (4.5, 4.6). While a general solution for the metric function is possible to obtain, we prefer not to present it here due to its cumbersome form. As in 4d [16], by choosing a particular relation between the hair parameters, i.e. $q_{c}=q_{b}/\kappa$, we can find a simpler solution given by
$f_{\pm}=-\frac{r^{2}}{2(\kappa-1)\ell^{2}}\left(1\pm\sqrt{1-4\Lambda_{0}(\kappa-1)\ell^{2}+\frac{4(\kappa-1)^{2}\ell^{4}(4q_{b}^{2}-\kappa mr^{2})}{\kappa r^{4}}}\,\right).$ (4.16)
The sign in front of the square root can be fixed by demanding a well defined $\ell\to 0$ limit. We have
$f_{+}$ $=-\frac{r^{2}}{(\kappa-1)\ell^{2}}+r^{2}\Lambda_{0}+\mathcal{O}(\ell^{2}),$ (4.17)
$f_{-}$ $=-r^{2}\Lambda_{0}+\mathcal{O}(\ell^{2}).$ (4.18)
which means that $f_{-}$ branch should be chosen. The effective cosmological constant can be read from the $r\to\infty$ limit as
$\Lambda=\frac{1-\sqrt{1-4(\kappa-1)\ell^{2}\Lambda_{0}}}{2(\kappa-1)\ell^{2}}.$ (4.19)
For an AdS3 black hole ($\Lambda<0$), $\Lambda_{0}$ should be negative. When $\kappa>1$, the solution is guaranteed to be asymptotically AdS3. However, when $\kappa<1$, the following additional constraint should be satisfied
$\Lambda_{0}\geq\frac{1}{4(\kappa-1)\ell^{2}}.$ (4.20)
As $r\to 0$, the Ricci scalar (and other curvature invariants) remain finite. Therefore, this is again a regular black hole solution.
4.4 Deformed theory ($\kappa\neq 1$) coupled to Born–Infeld electrodynamics
By the same procedure applied in Subsection 4.2, the metric function is obtained as
$f$ $=-\frac{r^{2}}{2(\kappa-1)\ell^{2}}\times$
$\left[1-\sqrt{1-\frac{2(\kappa-1)\ell^{2}(q_{e}^{2}+b^{2}r^{2})(b^{2}(\psi-1)+2\Lambda_{0})}{b^{2}r^{2}}-\frac{2(\kappa-1)\ell^{2}q_{e}^{2}\text{arctanh}(\psi)}{r^{2}}+\frac{4(\kappa-1)^{2}\ell^{4}(4q_{b}^{2}-r^{2}\kappa m)}{\kappa r^{4}}}\,\right]$ (4.21)
The effective cosmological constant $\Lambda$ is the same as in (4.19). It is straightforward to show that this is also a regular black hole solution.
5 Conclusion
In this paper, following [37], we obtained a 3d bi-vector-tensor theory by regularizing the GB invariant with two Weyl vectors and showed that it admits a regular AdS3 black hole solution with two distinct primary hair parameters. By introducing a deformation and choosing a particular relation between hair parameters, we obtain a different regular AdS3 black hole. Coupling the theories to Born-Infeld electrodynamics yields regular charged black holes.
With these solutions, it is possible to ask questions about the black hole interior for a regular black hole in 3d. Previous studies show that regular AdS3 black holes are geodesically complete and tidal forces experienced by observers are finite everywhere [13, 38]. The same behavior is expected for the solutions presented here. Because of the hair parameters, the thermodynamical properties of the black hole solutions given here might be particularly interesting (see [50, 51] for previous 3d examples).
The microscopic derivation of entropy by using the AdS3/CFT2 correspondence was generalized to solutions with scalar hair in [52, 53, 54]. Instead of the usual form of the Cardy formula for 2d CFTs with the central charge of the theory, an alternative version with the ground state energy was shown to successfully reproduce the semi-classical entropy of black holes with scalar hair. The ground state is identified on the gravitational side with the soliton solution obtained by a double Wick rotation from the static black hole solution. Together with the solutions in [14], we expect this procedure to work for solutions given here and we hope to report more on this in near future.
Appendix A Coupling to Maxwell’s electrodynamics
In this appendix, we give the resulting charged black hole solutions when the undeformed ($\kappa=1$) and the deformed ($\kappa\neq 1$) theory are coupled to Maxwell’s electrodynamics with the Lagrangian
$\mathcal{L}_{\text{M}}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}.$ (A.1)
Taking the vector field as in (4.11), one finds the Coulomb potential
$\phi=-q_{e}\log(r).$ (A.2)
The components of the Weyl vectors are given in (4.5, 4.6), and the metric functions are as follows
$\kappa=1:\quad\quad f$ $=\frac{r^{2}\left[-\Lambda_{0}r^{2}-m-\frac{q_{e}^{2}}{2}\log(r)\right]+4\ell^{2}(q_{b}^{2}-q_{c}^{2})}{r^{2}+4\ell^{2}(q_{b}-q_{c})},$ (A.3)
$\kappa\neq 1:\quad\quad f$ $=-\frac{r^{2}}{2\ell^{2}(\kappa-1)}\times$
$\left(1-\sqrt{1-4\Lambda_{0}(\kappa-1)\ell^{2}+\frac{4(\kappa-1)^{2}\ell^{4}(4q_{b}^{2}-\kappa mr^{2})}{\kappa r^{4}}-\frac{2(\kappa-1)q_{e}^{2}\ell^{2}\log(r)}{r^{2}}}\,\right).$ (A.4)
From the $r\to 0$ limit, the behavior of the Ricci scalar is found as
$\kappa=1:\quad\quad R$ $=\frac{12(q_{b}+q_{c}+m)+5q_{e}^{2}+6q_{e}^{2}\log(r)}{8\ell^{2}(q_{b}-q_{c})}+\mathcal{O}(r^{2}),$ (A.5)
$\kappa\neq 1:\quad\quad R$ $=\frac{12\kappa(\kappa-1)m\ell^{2}+5\kappa q_{e}^{2}+24\sqrt{\kappa}q_{b}+6\kappa q_{e}^{2}\log(r)}{8(\kappa-1)\sqrt{\kappa}\,\ell^{2}q_{b}}+\mathcal{O}(r^{2}).$ (A.6)
We see that these solutions are singular at the origin.